OpenStudy (anonymous):

evaluate (a-4b)^2 when a=15 and b=3 8th grade math PLEASE!!! please help very urgent

OpenStudy (anonymous):

\[(a-4b)^2 \]

OpenStudy (anonymous):

33?

OpenStudy (anonymous):

@lilprincess ??? this is more reasonable XD

OpenStudy (anonymous):

784

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

oh crap its \[(a-4b)^2+2a\]

OpenStudy (anonymous):

@perl mistake i made

OpenStudy (anonymous):

@charlotte123 ???

OpenStudy (charlotte123):

(a-4b)^2 a=15 and b=3 so (a-4b)^2 becomes (15-4*3)^2 ---> (15-12)^2

OpenStudy (charlotte123):

Your turn ~ ^_^

OpenStudy (anonymous):

k.

OpenStudy (anonymous):

i just started this subject soo...

OpenStudy (anonymous):

( ͡° ͜ʖ ͡°)

OpenStudy (charlotte123):

That face wow

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

i have options and i have to help my pissy brother very soon... can you help me out?

OpenStudy (charlotte123):

I...Idekrn That face...LEL What's 15 - 12?

OpenStudy (charlotte123):

Of course I can ^w^

OpenStudy (anonymous):

he already stabbed me with a pen :|

OpenStudy (charlotte123):

wow

OpenStudy (anonymous):

the options are as following

OpenStudy (charlotte123):

@crashcrow What's 15 - 12?

OpenStudy (anonymous):

3

OpenStudy (charlotte123):

Good! ^w^ And what's 3^2?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

heheh

OpenStudy (charlotte123):

And there's your answer ^w^ hehehe

OpenStudy (charlotte123):

Gj :')

OpenStudy (anonymous):

9?

OpenStudy (charlotte123):

mhmmm yep ^_^

OpenStudy (anonymous):

options are 39, 33, 11 and 36

OpenStudy (anonymous):

tho

OpenStudy (charlotte123):

a-4b)^2 a=15 and b=3 so (a-4b)^2 becomes (15-4*3)^2 ---> (15-12)^2 15 - 12 = 3 (3)^2 ^ = 9

OpenStudy (charlotte123):

wow....

OpenStudy (anonymous):

:/

OpenStudy (charlotte123):

that's real strange O_o

OpenStudy (anonymous):

yeah... ill have to email my teacher l8r about it

OpenStudy (charlotte123):

You should...that's just really strange ~

OpenStudy (anonymous):

ok ill go for 39 then ( ͡° ͜ʖ ͡°)

OpenStudy (charlotte123):

Hold on

OpenStudy (anonymous):

ok.

OpenStudy (charlotte123):

XD wow XD Are you supposed to find what's in the parenthesis or the overall answer?

OpenStudy (anonymous):

overall i guess. im soo lost @_@

OpenStudy (charlotte123):

So the answer should be 9 >_> (a-4b)^2 a=15 and b=3 so (a-4b)^2 becomes (15-4*3)^2 If you were to subtract 4 from 15 and then * by 3 you'd get 33...hmm...

OpenStudy (anonymous):

wow. okay ill try 33 and let you know. >.<

OpenStudy (charlotte123):

But then you'd have to multiply 33 by 33 because of the "^2"

OpenStudy (charlotte123):

Good luck >,>

OpenStudy (anonymous):

ill need it (>'.'>)

OpenStudy (charlotte123):

OH HOLD ON

OpenStudy (anonymous):

ok...

OpenStudy (charlotte123):

I didn't see where you corrected your mistake >,> It's really (a−4b)^2+2a =,=

OpenStudy (anonymous):

hm. 0-o

OpenStudy (charlotte123):

(a−4b)^2+2a a=15 and b=3 so (a−4b)^2+2a becomes (15-4*3)^2+2*15

OpenStudy (anonymous):

mmk

OpenStudy (charlotte123):

(15-4*3)^2+2*15 ----> 9 + 30

OpenStudy (charlotte123):

So you're answer's 39 ^w^

OpenStudy (anonymous):

39? kool thanks for the time

OpenStudy (charlotte123):

Mhmm ^w^ You're very welcome!

OpenStudy (anonymous):

ahh sh!t i got two more >.<

OpenStudy (charlotte123):

ohhhh wow I got work to do x_x

OpenStudy (anonymous):

np thank you much

OpenStudy (charlotte123):

You know what...go ahead ^_^

OpenStudy (anonymous):

so wow much thanks such mathing

OpenStudy (anonymous):

heheh

OpenStudy (anonymous):

ill make a new post

OpenStudy (charlotte123):

LOL You know how to speak doge :')

OpenStudy (charlotte123):

Okay XD